### v1

# The Experimental Unit Information Index: Balancing Evidentiary Value and Sample Size of Adaptive Designs

Abstract:
Reducing the number of experimental units is one of the three pillars of the 3R principles (Replace, Reduce, Refine) in animal research. At the same time, statistical error rates need to be controlled to enable reliable inferences and decisions. This paper proposes a novel measure to quantify the evidentiary value of one experimental unit for a given study design. The experimental unit information index (EUII) is based on power, Type-I error and sample size, and has attractive interpretations both in terms of frequentist error rates and Bayesian posterior odds. We introduce the EUII in simple statistical test settings and show that its asymptotic value depends only on the assumed relative effect size under the alternative. We then extend the definition to adaptive designs where early stopping for efficacy or futility may cause reductions in sample size. Applications to group-sequential designs and a recently proposed adaptive statistical test procedure show the usefulness of the approach when the goal is to maximize the evidentiary value of one experimental unit.
Key Words: Adaptive designs; Diagnostic odds ratio; Likelihood ratio; 3R Principles; Bayesian updating
1 Introduction
Standard statistical hypothesis testing is based on the control of the Type-I error rate at some pre-defined significance level $\alpha$. “Control” means that the actual Type-I error rate is not allowed to be larger than $\alpha$, but it can be smaller. A comparison of different statistical tests that control the Type-I error rate is then based on the power to detect a certain effect size of interest. The greater the power, the better the test. However, power comparisons are rarely meaningful if the Type-I error rates of the tests are different, as a greater power can typically be traded-off with an increase in Type-I error rate.
Over the last decades, various adaptive designs have been developed where an experiment can be stopped early, either for efficacy (a convincing result has already been obtained) or for futility (the results are so unconvincing that there is only little sense in continuing the experiment). Stopping an experiment early ensures that the number of animals included in an experiment does not become too large. This reflects the second pillar of the 3R principles “Replace, Reduce, Refine” in animal research (Russell and Burch, 1959), described as
to use methods for obtaining comparable levels of information from the use of fewer animals in scientific procedures, or for obtaining more information from the same number of animals” (swiss3rcc.org/3rs-for-the-public/)
The 3R principles are now widely adopted to improve the design and analysis of animal studies (Reynolds, 2023, 2024a, 2024b; Townsend et al., 2025). However, the focus of standard statistical theory on error rates ignores that different adaptive designs may have different (expected) sample sizes. Furthermore, a comparison of methods with desired Type-I error rate control with alternative methods that may have more power but also Type-I error rate inflation can be difficult. To address the latter issue, optimising the trade-off between Type-I and Type-II error rates has been recently considered Mudge et al. (2012); Walley and Grieve (2021); Grieve (2024). The goal is here to optimize a weighted sum of error rates based on prior probabilities of the null and the alternative hypothesis and the relative costs of Type-I and Type-II errors. An alternative approach is to adjust power for the corresponding Type-I error rate using the theory of receiver operating characteristic curves (Lloyd, 2005; Cavus et al., 2019). However, all these methods do not take sample size into account.
In this paper we propose a new measure of statistical information that takes into account Type-I error rate, power, and sample size. This measure is related to the diagnostic odds ratio to quantify diagnostic accuracy (Glas et al., 2003), which is widely used in clinical epidemiology and which also has attractive interpretations when applied to statistical hypothesis tests. First, it has a frequentist interpretation as the ratio of the rejection odds under the alternative and under the null hypothesis. Second, it also has a Bayesian interpretation as the ratio of the posterior odds of the alternative hypothesis given a significant versus nonsignificant results. Finally, the diagnostic odds ratio allows for a normalization with respect to the (expected) sample size to obtain a measure of the evidentiary value of one unit/animal/patient, the experimental unit information index (EUII).
This article is organized as follows: Section 2 reviews likelihood ratios and diagnostic odds ratios and applies them to statistical hypothesis tests. Section 3 then defines the experimental unit information index for fixed sample size and studies its performance in standard statistical test settings. The extension to adaptive tests with different expected sample size of significant and nonsignificant findings is described in Section 4. Applications to group-sequential designs (Section 4.1) and “constrained sample augmentation” (Section 4.2), an adaptive test procedure recently proposed by Reinagel (2023), are described. We close with some discussion in Section 5.
2 Methodology
2.1 Likelihood ratios and the diagnostic odds ratio
Sensitivity and specificity are traditional measures of diagnostic test accuracy (Pepe, 2004). Sensitivity is defined as the probability to obtain a positive test result from a diseased patient. Specificity is defined as the probability of a negative test result from a healthy patient.
In a diagnostic context, $H_{1}$ represents the hypothesis that a patient has the disease whereas $H_{0}$ represents that the patient does not have the disease. Likelihood ratios (LRs) are commonly used in diagnostic testing to quantify the (prior to posterior) change in odds $\mathsf{Odds}(H_{1})=\Pr(H_{1})/\Pr(H_{0})$ after having observed a positive or negative test result (Deeks and Altman, 2004):
$\left.\begin{array}[]{c}{\mathsf{Odds}(H_{1}\,|\,\mbox{test positive})}\\ {\mathsf{Odds}(H_{1}\,|\,\mbox{test negative})}\end{array}\right\}$ $=$ $\left\{\begin{array}[]{c}\mbox{LR}^{+}\\ \mbox{LR}^{-}\end{array}\right\}\,\cdot\,{\mathsf{Odds}(H_{1})}.$ (5)
LRs are useful because they do not depend on the pre-test odds of having the disease (typically the prevalence), which may be unknown or differ across populations.
Positive and negative likelihood ratios directly relate to sensitivity and specificity,
positive LR: $\mbox{LR}^{+}=$ $\frac{\Pr(\text{test positive}\,|\,\text{diseased})}{\Pr(\text{test positive}\,|\,\text{non-diseased})}=\frac{\mbox{Sensitivity}}{1-\mbox{Specificity}}$
negative LR: $\mbox{LR}^{-}=$ $\frac{\Pr(\text{test negative}\,|\,\text{diseased})}{\Pr(\text{test negative}\,|\,\text{non-diseased})}=\frac{1-\mbox{Sensitivity}}{\mbox{Specificity}}$
and the diagnostic odds ratio (Glas et al., 2003) (DOR) is simply the ratio of the two likelihood ratios: $\mbox{DOR}=\mbox{LR}^{+}/\mbox{LR}^{-}$.
Useful diagnostic tests have positive likelihood ratio $\mbox{LR}^{+}>1$ and negative likelihood ratio $\mbox{LR}^{-}<1$, where one of the two inequalities implies the other one. Equation (5) shows that a positive result from a useful test increases the odds (and therefore also the probability) of having the disease whereas a negative result decreases it. The diagnostic odds ratio is thus a single indicator of diagnostic test accuracy, the larger, the better. We will describe its properties in more detail in the following, when we apply it to statistical hypotheses tests.
2.2 The diagnostic odds ratio for hypotheses testing
It has been argued that likelihood ratios can also be applied to statistical hypothesis tests (Bayarri et al., 2016; Huang and Trinquart, 2024), where a positive or negative diagnostic test result is now replaced by a significant or nonsignificant hypothesis test result. The positive and negative likelihood ratios are then known as the (pre-experimental) rejection ratio (Bayarri et al., 2016) and nonrejection ratio (Phelan, 2018), respectively. Huang and Trinquart (2024) apply the positive and negative likelihood ratio to hypothesis tests and consider
positive LR: $\mbox{LR}^{+}=$ $\frac{\Pr(\mbox{significant}\,|\,H_{1})}{\Pr(\mbox{significant}\,|\,H_{0})}=\frac{\mbox{Power}}{\mbox{T1E rate}},$ (6)
negative LR: $\mbox{LR}^{-}=$ $\frac{\Pr(\mbox{nonsignificant}\,|\,H_{1})}{\Pr(\mbox{nonsignificant}\,|\,H_{0})}=\frac{1-\mbox{Power}}{1-\mbox{T1E rate}}.$ (7)
For a significant finding, the positive likelihood ratio directly relates to the positive predictive value $\Pr(H_{1}\,|\,\mbox{significant})$ (Browner, 1987) via
$\mathsf{Odds}(H_{1}\,|\,\mbox{significant})=\mbox{LR}^{+}\cdot\mathsf{Odds}(H_{1}).$ (8)
Similarly, the negative likelihood ratio $\mbox{LR}^{-}$ directly relates to the negative predictive value $\Pr(H_{0}\,|\,\mbox{nonsignificant})=1-\Pr(H_{1}\,|\,\mbox{nonsignificant})$:
$\mathsf{Odds}(H_{1}\,|\,\mbox{nonsignificant})=\mbox{LR}^{-}\cdot\mathsf{Odds}(H_{1}).$ (9)
As for diagnostic tests, useful statistical tests have positive likelihood ratio $\mbox{LR}^{+}>1$ and negative likelihood ratio $\mbox{LR}^{-}<1$. This means that a significant test result increases the odds (and the probability) of the alternative hypothesis $H_{1}$ being true, whereas a nonsignificant result increases the odds of the null hypothesis $H_{0}$. In what follows we describe the approach for standard statistical hypothesis tests but note that the approach can also be applied to other inferential approaches such as dual-criterion designs (Roychoudhury et al., 2018) or Bayesian methods (Spiegelhalter et al., 2004), as long as these methods provide a dichotomous test result in the form of accepting either $H_{1}$ (“significant”) or not (“not significant”).
The diagnostic odds ratio of a statistical test can now be defined as the ratio of the positive likelihood ratio (6) and the negative likelihood ratio (7):
$\mbox{DOR}=\frac{\mbox{LR}^{+}}{\mbox{LR}^{-}}=\frac{{\mbox{Power}}/{\mbox{T1E rate}}}{(1-{\mbox{Power}})/(1-{\mbox{T1E rate}})}.$ (10)
Rearranging this equation gives the frequentist interpretation of the DOR as the ratio of the odds of significance under $H_{1}$ (“Power Odds”) to the odds of significance under $H_{0}$ (“T1E Odds”):
DOR $=$ $\frac{\mbox{Power Odds}}{\mbox{T1E Odds}}$ (11)
where
$\mbox{Power Odds}=\frac{\mbox{Power}}{1-\mbox{Power}}$ and $\mbox{T1E Odds}=\frac{\mbox{T1E rate}}{1-\mbox{T1E rate}}.$
DOR thus describes the evidentiary value of a statistical test in discriminating between $H_{1}$ and $H_{0}$. The DOR ranges from 0 to infinity and is larger than 1 for a useful statistical test. If $\mbox{DOR}=1$, the power is equal to the T1E rate, so the probability of significance is the same under $H_{1}$ and $H_{0}$ and the test is useless.
The DOR can also be written as the ratio of the odds of $H_{1}$ given a significant result relative to the odds of $H_{1}$ given a nonsignificant result:
DOR $=$ $\frac{\mathsf{Odds}(H_{1}\,|\,\mbox{significant})}{\mathsf{Odds}(H_{1}\,|\,\mbox{nonsignificant})}$ (12)
This formulation directly follows from division of (8) with (9) and is particularly remarkable, because it is a statement about posterior odds of the alternative $H_{1}$ without specification of the prior probability $\Pr(H_{1})$ or the corresponding prior odds (the prior odds $\mathsf{Odds}(H_{1})$ cancel in the division of (8) with (9)). So the DOR quantifies how much larger the odds of the alternative $H_{1}$ are among significant results compared to nonsignificant results. A similar rearrangement of Bayes’ theorem led to the crucial understanding that the odds ratio in a retrospective case-control study can be interpreted prospectively (Cornfield, 1951; Breslow, 1996). Note that (12) can also be formulated in terms of odds of the null hypothesis $H_{0}$:
DOR $=$ $\frac{\mathsf{Odds}(H_{0}\,|\,\mbox{nonsignificant})}{\mathsf{Odds}(H_{0}\,|\,\mbox{significant})}.$ (13)
To summarize, the DOR quantifies how large the odds of $H_{1}$ of significant test results are compared to nonsignificant test results. Equivalently, DOR tells us how large the odds of $H_{0}$ of nonsignificant test results are compared to significant test results. The DOR has therefore useful interpretations both in terms of frequentist error rates (11) and Bayesian posterior odds (12) and (13).
3 The experimental unit information index
To adjust the DOR for a pre-specified sample size $n$, we propose the experimental unit information index as the $n$th root of the diagnostic odds ratio:
$\mbox{EUII}=\frac{\left(\mbox{LR}^{+}\right)^{1/n}}{\left(\mbox{LR}^{-}\right)^{1/n}}=\mbox{DOR}^{1/n}.$ (14)
The motivation to take the $n$th root comes from equation (8) and (9), respectively. These two equations show that likelihood ratios act multiplicatively on the odds of $H_{1}$, transforming prior odds to posterior odds. Both likelihood ratios are based on data from $n$ experimental units, so evey unit contributes on average an amount of $({\mbox{LR}^{-}})^{1/n}$ and $({\mbox{LR}^{+}})^{1/n}$, respectively, to the likelihood ratios, because
$\mbox{LR}=\underbrace{{\mbox{LR}}^{1/n}\times{\mbox{LR}}^{1/n}\times\ldots\times{\mbox{LR}}^{1/n}}_{\mbox{$n$ times}}.$
This representation emphasizes the coherence of Bayesian (prior to posterior) updating based on increasingly available data. Taking the ratio of $({\mbox{LR}^{-}})^{1/n}$ and $({\mbox{LR}^{+}})^{1/n}$ leads to the definition (14) of the experimental unit information index (EUII). Note that this derivation assumes that the experimental units are independent. The EUII hence represents the average contribution of one experimental unit to the DOR of a statistical test, where “average” is to be understood as the geometric average. A useful statistical test has EUII > 1, so on average each additional experimental unit provides information that helps to discriminate between the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{1}$).
Taking the logarithm of (14),
$\log\mbox{EUII}=\frac{1}{n}\log\mbox{DOR},$
shows that the log EUII can be interpreted as the average contribution of each observation to the log diagnostic odds ratio. Further, writing the $\log\mbox{EUII}$ as
$\log\mbox{EUII}=\frac{1}{n}\left\{\log\left(\frac{1}{1-\mbox{Power}}-1\right)+\log\left(\frac{1}{\mbox{T1E rate}}-1\right)\right\}$
shows explicitly that two tests can have the same EUII for different combinations of power, T1E rate, and sample size. This means that if less than two out of three quantities are fixed, there is possibly no unique statistical test that has the highest EUII. Fixing the sample size and the T1E rate results in the classical uniformly most powerful test emerging as the test with the highest EUII. Conversely, fixing the power and T1E rate, the test with the lowest sample size emerges as the optimal test that maximizes the EUII. Finally, fixing the sample size and the power, the test with the lowest T1E rate is the optimal test that maximizes the EUII. All of these three cases are in line with common sense.
3.1 Empirical evaluation
We will now evaluate the EUII for both $z$ and $t$-tests.
3.1.1 Standard z-test
For simplicity we consider a standard two-sided $z$-test based on one sample, where there is a closed-form expression for the required sample size to achieve a power $1-\beta$ to detect a relative effect size $\delta$ (Kirkwood and Sterne, 2003)
$n=\frac{(u+v)^{2}}{\delta^{2}}$ (15)
with Type-II error (T2E) rate $\beta$, $u=\Phi^{-1}(1-\beta)$ and $v=\Phi^{-1}(1-\alpha/2)$. The required sample size $n$ is usually rounded up to the next integer and then the actual power will be slightly larger than $1-\beta$. However, we will work in the following with the unrounded sample size from (15) to keep the power exactly at the desired value.
For example, suppose we design a study to achieve a power of $1-\beta=80\%$ to detect the relative effect size $\delta=0.5$ at the (two-sided) significance level of $\alpha=5\%$. We then have $\mbox{LR}^{+}={\mbox{80\%}}/{\mbox{5\%}}=16$, $\mbox{LR}^{-}={\mbox{20\%}}/{\mbox{95\%}}=4/19$ and so $\mbox{DOR}=16/(4/19)=76$. Alternatively, the DOR can be computed as the ratio of the $\mbox{Power Odds}={\mbox{80\%}}/{\mbox{20\%}}=4$ and the $\mbox{T1E Odds}={\mbox{5\%}}/{\mbox{95\%}}=1/19$, which then also gives $\mbox{DOR}=4\cdot 19=76$. The required sample size (15) to achieve 80% power to detect $\delta=0.5$ with a one-sample $t$-test is $n=31.4$. The EUII is therefore
$\mbox{EUII}_{1}=76^{1/31.4}=1.15,$
which means that every experimental unit in the sample increases the evidentiary value of the statistical test on average by 15%.
Table 1 gives the DOR and the EUII also for other values of the power and Type-I error rate. We can see that a test with 90% power at significance level 5% has more evidentiary value than a test with 80% power at significance level 2.5% (DOR = 171 vs. 156). The increase in DOR is parallelized by an increase in the required sample size $n$ to detect the standardized effect $\delta=0.5$, so the corresponding values of EUII are nearly the same, 1.13 and 1.14, respectively. The EUII takes its largest value 1.15 if both Type-I and Type-II error have the largest value (5% and 20%, respectively). We will see later, that the asymptotic value of EUII (for fixed T1E rate and increasing sample size $n$) is $\exp(\delta^{2}/2)=1.133$, so the finite sample size values of EUII in Table 1 are all close to this asymptotic value.
Table 1: Diagnostic odds ratio (DOR) and experimental unit information index (EUII) for different values of power and Type-I error rate. Also shown is the required sample size $n$ to detect a standardized effect of $\delta=0.5$ with a one-sample $z$-test and the corresponding value of EUII.
Power (in %) Type-I error rate (in %) DOR $n$ EUII
80 5.0 76.0 31.4 1.15
80 2.5 156.0 38.0 1.14
90 5.0 171.0 42.0 1.13
90 2.5 351.0 49.6 1.13
If the design would have been instead based on two equally-sized samples and $\delta$ represents the standardized difference to be detected, then the required total sample size $n=125.6$ at the 5% significance level and 80% power is four times as large (Kirkwood and Sterne, 2003), and we obtain
$\mbox{EUII}_{2}=76^{1/125.6}=1.035.$
The average evidentiary increase per unit is now 3.5%. Note that $\mbox{EUII}_{1}=({\mbox{EUII}_{2}})^{4}$ which means that one unit in a one-sample study contributes as much information as four units in a (perfectly balanced) two-sample trial, if the goal is to detect the same standardized effect.
Note that we take $n$ to be the total sample size and not the sample size per group. This enables us to adjust EUII for unequal sample sizes. Unequal sample sizes are quite common in clinical trials even if allocation is based on simple 1:1 randomisation. Furthermore, unequal allocation (e.g., 2:1) can provide investigators with greater experience of a new treatment and may even encourage recruitment to the trial.
The effect of unequal sample sizes on the power of the $z$-test has been discussed in Matthews (2006). If we have $n=126$ patients in total, the power will decrease to
$1-\beta=\Phi\left(\delta\sqrt{\frac{n_{1}\,n_{2}}{126}}-v\right)$
where $n_{1}$ patients are allocated to one group and $n_{2}=126-n_{1}$ to the other. Unequal sample sizes will thus decrease the power odds, the DOR and hence also the EUII, because the total sample size $n=126$ remains fixed. This is illustrated in Figure 1.
Figure 1: The experimental unit information index for unequal randomisation of $n=126$ patients to two groups. The red line at $n=42$ corresponds to a 2:1 randomisation.
For example, if the sample size in the smaller group is $n_{1}=42$ (corresponding to a 2:1 allocation ratio), then EUII reduces from 1.035 to slightly below 1.033. This confirms that the information loss of unequal randomisation in terms of EUII is small if the allocation ratio is not greater than 2:1 (Matthews, 2006).
3.2 Z-test and t-test with increasing sample size
Consider now the standard one-sample one-sided $t$-test to test the null hypothesis $H_{0}\colon\mu=0$ versus the alternative $H_{1}\colon\mu>0$. We let the sample size $n$ vary between $2^{3}=8$ and $2^{13}=8192$ and compute the power to detect a standardized effect size $\delta$ at significance level $\alpha=5\%$. Figure 2 shows how the resulting EUII varies with $\delta$ and $n$. Not surprisingly, EUII increases with increasing $\delta$, because the power to detect $\delta$ increases with increasing $\delta$ (for fixed $n$). If $\delta=0.1$ and $n=8$, EUII is smaller than 1 because the power to detect such a small effect with such a small sample size is in fact smaller than the T1E rate. Such a test is called biased in the mathematical statistics literature (Lehmann, 1959).
Figure 2: The experimental unit information index for the standard one-sample one-sided $t$-and $z$-test as a function of sample size and effect size. The asymptotic value $\exp(\delta^{2}/2)$ for the different values of the standardized effect size $\delta$ is indicated by the dotted lines.
Note that the power also increases with increasing $n$, but the EUII index adjusts for the sample size $n$ and is able to compensate for this increase. It first drops and then increases to the finite limiting value
$\lim_{n\rightarrow\infty}\mbox{EUII}_{1}=\exp\left(\frac{\delta^{2}}{2}\right)$ (16)
as shown in Appendix A. For comparison, Figure 2 also shows EUII for the corresponding $z$-test, which is slightly larger for small $n$, but converges to the same limit.
In the two-sample case with equal group sizes, the limiting value of EUII is the fourth root of (16), i. e.
$\lim_{n\rightarrow\infty}\mbox{EUII}_{2}=\exp\left(\frac{\delta^{2}}{8}\right).$ (17)
Thus the asymptotic evidentiary gain from one additional unit is lower for a two-sample compared to a one-sample test. Some rearrangement of (16) and (17) shows that the effect size $\delta$ in the two-sample setting needs to be twice as large as in the one-sample setting to obtain the same asymptotic evidentiary value of one experimental unit.
4 The experimental unit information index for adaptive designs
In adaptive trial designs, a trial can often be stopped preliminary if the results are sufficiently compelling at an interim analysis, for example, if there is strong evidence for treatment benefit or harm (Kairalla et al., 2012). This means that the sample size is random and typically differs under $H_{0}$ and $H_{1}$. In the following, we denote a random sample size by $N$ to distinguish it from a fixed sample size $n$. An adaptive trial can then be characterized by the pair of the sample size $N$ at which the trial stops and the indicator $S$ of whether the results were (non)significant. The expected sample sizes for significant (++) and nonsignificant (-) are then then defined by the conditional expectations $\operatorname{\mathsf{E}}(N_{+})=\operatorname{\mathsf{E}}(N\,|\,S=\text{significant})$ and $\operatorname{\mathsf{E}}(N_{-})=\operatorname{\mathsf{E}}(N\,|\,S=\text{nonsignificant})$, respectively.
Differences in sample sizes make it difficult to compare operating characteristics (such as power and T1E rate) of different trial designs. The fixed sample size definition (14) suggests a way to generalize the EUII to settings with random sample size:
$\mbox{EUII}=\frac{\big(\mbox{LR}^{+}\big)^{1/\operatorname{\mathsf{E}}(N_{+})}}{\big(\mbox{LR}^{-}\big)^{1/\operatorname{\mathsf{E}}(N_{-})}}.$ (18)
The rationale for the generalization (18) is that the computation of posterior predictive odds with likelihood ratios is either based on significant (equation (8)) or nonsignificant findings (equation (9)). It is therefore natural to normalize the numerator $\mbox{LR}^{+}$ in (18) with the expected sample size of significant findings $\operatorname{\mathsf{E}}(N_{+})$ and the denominator with the expected sample size of nonsignificant findings $\operatorname{\mathsf{E}}(N_{-})$.
For designs with fixed sample size $n=\operatorname{\mathsf{E}}(N_{+})=\operatorname{\mathsf{E}}(N_{-})$, this generalization reduces correctly to (14). Mathematically, equation (18) can be justified by a first-order Taylor approximation of the terms
[…]
in the expected log EUII
$\operatorname{\mathsf{E}}(\log\mbox{EUII})=\operatorname{\mathsf{E}}(1/N_{+})\log\mbox{LR}^{+}-\operatorname{\mathsf{E}}(1/N_{-})\log\mbox{LR}^{-},$
where $\log\mbox{LR}^{+}$ and $\log\mbox{LR}^{-}$ are assumed to be fixed. However, a first-order Taylor approximation may be inaccurate. The second order Taylor approximations
[…]
also incorporates the coefficient of variation
[…]
respectively, and hence the variance of the sample size, leading to the second-order approximation
$\widetilde{\mbox{EUII}}=\frac{\left(\mbox{LR}^{+}\right)^{\left(1+\mathsf{CV}(N_{+})^{2}\right)/\operatorname{\mathsf{E}}(N_{+})}}{\big(\mbox{LR}^{-}\big)^{\left(1+\mathsf{CV}(N_{-})^{2}\right)/\operatorname{\mathsf{E}}(N_{-})}}.$ (19)
One can view $\operatorname{\mathsf{E}}(N_{\bullet})/(1+\mathsf{CV}(N_{\bullet})^{2})$ (where $\bullet\in\{+,-\}$) as an adjusted expected sample size that also takes the coefficient of variation into account. The larger the coefficient of variation, the smaller the adjusted expected sample size will be. The difference between (18) and (19) will be small if $\mathsf{CV}(N_{+})$ and $\mathsf{CV}(N_{+})$ are small. We will study the difference in the application described in Section 4.2.
A computational complication arises because significant and nonsignificant results can occur both under the null and alternative hypothesis. Monte Carlo simulation can often be used to obtain the distribution of the sample sizes $N_{0}$ and $N_{1}$ under the null and alternative hypothesis, respectively. We can then compute the expected sample size for significant results separately under the null ($\operatorname{\mathsf{E}}(N_{0+})$) and alternative ($\operatorname{\mathsf{E}}(N_{1+})$), and likewise for nonsignificant results ($\operatorname{\mathsf{E}}(N_{0-})$ and $\operatorname{\mathsf{E}}(N_{1-})$), see Table 2 for notation.
Table 2: The expected sample size under different hypotheses and test results.
Test result
Hypothesis significant nonsignificant
$H_{0}$ $\operatorname{\mathsf{E}}(N_{0+})$ $\operatorname{\mathsf{E}}(N_{0-})$ $\operatorname{\mathsf{E}}(N_{0})$
$H_{1}$ $\operatorname{\mathsf{E}}(N_{1+})$ $\operatorname{\mathsf{E}}(N_{1-})$ $\operatorname{\mathsf{E}}(N_{1})$
$\operatorname{\mathsf{E}}(N_{+})$ $\operatorname{\mathsf{E}}(N_{-})$ $\operatorname{\mathsf{E}}(N)$
But the expected sample size is in general different under the two alternatives, and then we have
$\operatorname{\mathsf{E}}(N_{+})$ $=$ $\operatorname{\mathsf{E}}(N_{0+})\Pr(H_{0}\,|\,\mbox{significant})+\operatorname{\mathsf{E}}(N_{1+})\Pr(H_{1}\,|\,\mbox{significant})$ (20)
$\operatorname{\mathsf{E}}(N_{-})$ $=$ $\operatorname{\mathsf{E}}(N_{0-})\Pr(H_{0}\,|\,\mbox{nonsignificant})+\operatorname{\mathsf{E}}(N_{1-})\Pr(H_{1}\,|\,\mbox{nonsignificant})$ (21)
for significant and nonsignificant results, respectively. The probability $\Pr(H_{1}\,|\,\mbox{significant})$ depends on the prior probability $\Pr(H_{1})$ through (8), and is usually unknown. All other probabilities in (20) and (21) are functions of $\Pr(H_{1}\,|\,\mbox{significant})$:
$\Pr(H_{0}\,|\,\mbox{significant})$ $=$ $1-\Pr(H_{1}\,|\,\mbox{significant})$
$\Pr(H_{1}\,|\,\mbox{nonsignificant})$ $=$ $\left(1+\left[\mbox{DOR}^{-1}\cdot\frac{\Pr(H_{1}\,|\,\mbox{significant})}{1-\Pr(H_{1}\,|\,\mbox{significant})}\right]^{-1}\right)^{-1}$
$\Pr(H_{0}\,|\,\mbox{nonsignificant})$ $=$ $1-\Pr(H_{1}\,|\,\mbox{nonsignificant}),$
where the second equation follows from (12).
To compute the second-order approximation (19), we also need $\operatorname{Var}(N_{+})$ and $\operatorname{Var}(N_{-})$. The law of total probability for variances (De Groot and Schervish, 2012) gives
$\operatorname{Var}(N_{+})$ $=$ $\operatorname{Var}(N_{0+})\Pr(H_{0}\,|\,\mbox{significant})+\operatorname{Var}(N_{1+})\Pr(H_{1}\,|\,\mbox{significant})$ (22)
$+$ $\left[\operatorname{\mathsf{E}}(N_{0+})-\operatorname{\mathsf{E}}(N_{+})\right]^{2}\Pr(H_{0}\,|\,\mbox{significant})$
$+$ $\left[\operatorname{\mathsf{E}}(N_{1+})-\operatorname{\mathsf{E}}(N_{+})\right]^{2}\Pr(H_{1}\,|\,\mbox{significant})$
and likewise for $\operatorname{Var}(N_{-})$. Combined with (20) and (21) this gives $\mathsf{CV}(N_{+})$ and $\mathsf{CV}(N_{-})$ to be used in (19).
To summarize, in the case of varying sample size the EUII (18) depends on the expected sample size for significant and nonsignificant findings. A generalization also depends on the coefficient of variation. Both expectation and coefficient of variation depend on an additional input variable, the prior probability $\Pr(H_{1})$ (or $\Pr(H_{0})=1-\Pr(H_{1})$) of the alternative (null) hypothesis. Specification of this probability depends on the scientific context of the experiment. For example, in clinical research a number of 10% for the percentage of effective treatments has been suggested (Staquet et al., 1979; Simon, 1982). In discovery-oriented research this probability is expected to be even lower (Ioannidis, 2005).
4.1 Group-sequential designs
We now compare the EUII of the Pocock, O’Brien-Fleming and Haybittle-Peto group-sequential methods (Matthews, 2006) with four equally-spaced interim analyses (including the final analysis) in the one-sample case with the EUII from a fixed sample size design with $n=50$. Table 3 gives various characteristics of the different methods if we apply a $z$-test with known residual variance. Calculations are based on the gsDesign package in R. The package computes expected sample sizes under $H_{0}$ and $H_{1}$, which we have used to calculate $\operatorname{\mathsf{E}}(N_{+})$ and $\operatorname{\mathsf{E}}(N_{-})$. However, variances are not available so we have applied the first-order approximation (18) of EUII.
Table 3: Characteristics of the considered group-sequential designs. Shown are the nominal significance levels $a_{1},\dots,a_{4}$ at the four interim analyses, the Type-I error (T1E) rate, and the maximum possible sample size $n_{\scriptsize\texttt{Max}}$ to achieve a power of 90% as with a fixed design with $n=50$ and T1E rate of $2.5\%$. For Haybittle-Peto, $n_{\scriptsize\texttt{Max}}$ is used as for the fixed design.
Method $a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ T1E rate (in %) $n_{\scriptsize\texttt{Max}}$
Pocock 0.00911 0.0091 0.0091 0.0091 2.50 59.2
O’Brien-Fleming 0.00003 0.0021 0.0097 0.0215 2.50 51.1
Haybittle-Peto 0.00050 0.0005 0.0005 0.0250 2.54 50.0
Figure 3: The EUII (top) of different group-sequential methods with four interim analyses in comparison to a fixed design with $n=50$. The maximum sample size $n_{\scriptsize\texttt{Max}}$ of O’Brien-Fleming and Pocock is chosen so that their power (top axis) and Type-I error rate (2.5%) match with the fixed design, whereas $n_{\scriptsize\texttt{Max}}$ from Haybittle-Peto is always $n=50$, leading to slightly higher power and Type-I error rate. Middle: Improvement in EUII relative to the fixed design. Bottom: Maximum sample size $n_{\scriptsize\texttt{Max}}$ and expected sample size of significant findings. The prior probability $\Pr(H_{1})$ is assumed to be 0.01, 0.1, and 0.5
Note that Pocock and O’Brien-Fleming exactly control the Type-I error rate at $\alpha=2.5\%$ whereas the one-sided Haybittle-Peto method (Blenkinsop et al., 2019) (with nominal thresholds $a_{1}=a_{2}=a_{3}=0.0005$ and $a_{4}=0.025$) has a slightly increased overall T1E rate of 2.54%. The maximum sample size $n_{\scriptsize\texttt{Max}}$ (last column of Table 3) from Pocock and O’Brien-Fleming is larger than the corresponding sample size 50 under a fixed design with the same Type-I error rate and power of 90% (corresponding to a standardized effect size of $\delta=0.46$). Without futility stopping, this is also the sample size for any nonsignificant result, both under the null and the alternative: $\operatorname{\mathsf{E}}(N_{0-})=\operatorname{\mathsf{E}}(N_{1-})=\operatorname{\mathsf{E}}(N_{-})=n_{\scriptsize\texttt{Max}}$. However, the expected sample size for significant results will be smaller and in general depends on the underlying hypothesis and the assumed power. The maximum sample size $n_{\scriptsize\texttt{Max}}=50$ from Haybittle-Peto is chosen to be equal to the sample size from the fixed design. Although the power and Type-I error rate of Haybittle-Peto does not correspond in this case to the other methods, the EUII enables a meaningful comparison.
Figure 3 displays the EUII for the different methods (top), relative to the fixed design (middle), and the maximum sample size and expected sample size of significant findings (bottom) for values of the standardized effect size $\delta$ between 0.1 and 0.8. As expected, the EUII increases with increasing $\delta$ but also depends on the assumed prior probability of $H_{1}$, chosen to be either 0.01, 0.1 or 0.5. In particular, for $\Pr(H_{1})=.01$ the EUII depends less on the effect size $\delta$ under the alternative, because the second term $\operatorname{\mathsf{E}}(N_{1+})$ (that depends on $\delta$) of the expected sample size (20) for significant findings receives less weight.
All group-sequential designs have a larger EUII than the fixed design. The improvement is around 1-11% for the Pocock method, and around 0-7% for O’Brien-Fleming and Haybittle-Peto. These results indicate that Pocock should be the preferred group-sequential method if the goal is to maximize the evidentiary value of one experimental unit.
Figure 3 also shows the expected sample size of significant findings under the two hypotheses, for the different values of the standardized effect size. With the exception of Pocock, the expected sample size of significant findings is always larger under $H_{0}$ than under $H_{1}$. Only for Pocock it can be the other way round if the standardized effect size is less than $\delta=0.46$, which corresponds to the power being below 90%. This is a consequence of the fact that, under $H_{0}$, a significant finding with Pocock’s method is more likely to be obtained at the first interim analysis and less at the second and so forth. This is not the case under $H_{1}$, where a significant finding at the second interim analysis is more likely than at the first. This peculiar feature is caused by the constant and relatively large nominal significance levels (see Table 3) and has been one of the criticisms of the Pocock method in clinical trials, where – under the $H_{0}$ – it is less difficult that a trial is stopped early than for competing methods (Simon, 1994; Matthews, 2006). However, this feature has been described as an adventage in animal trials where the goal is to minimize animal numbers (Blotwijk et al., 2022) and our analysis confirms that smaller expected sample sizes under $H_{0}$ lead to a larger evidentiary value per animal of the Pocock method compared to the other ones. For example, the EUII is smaller for Haybittle-Peto, where the nominal significance levels (with the exception of the last one) are also constant, but smaller. Significant findings are then more likely to be obtained at the last analysis and the expected sample size under $H_{0}$ increases accordingly. In sum, the EUII allowed us to compare sequential designs with differing power, Type-I error rate, sample sizes, and interim analyses time points.
4.2 N-hacking and constrained sample augmentation
The term $N$-hacking was recently introduced by Reinagel (2023) to describe an adaptive procedure where more data is collected in an effort to achieve statistical significance, if only a “trend to significance” is observed so far. Without appropriate adjustment of the significance level $\alpha$ (usually 5% for a two-sided test), such a “peeking” procedure is known to inflate the Type-I error rate (Ludbrook, 2003). However, Reinagel (2023) argues that with an additional stop for futility if $p>0.1$, the Type-I error rate is only moderately increased while the positive predictive value $\Pr(H_{1}\,|\,\mbox{significant})$, the probability that a significant result is true positive, is substantially larger than for a fixed design with significance level $\alpha$. In addition, the mean final sample size is smaller compared to fixed sample size experiments with the same Type-I error rate and power. This aligns with the “Reduce” pillar of the 3R principles.
Reinagel specifically considers a modification of $N$-hacking, so-called “constrained sample augmentation”, to include an additional stop for futility: First, data are collected until 8 observations are collected, 4 in each group. If an independent sample Student’s $t$-test produces two-sided $p\leq 0.05$, significance is achieved, whereas $p>0.1$ leads to a futility stop of the experiment. If $0.05<p\leq 0.1$, another four observations are added and the updated $p$-value is used again to decide whether to stop for significance, stop for futility, or to continue with another four observations. This procedure is continued until a total of $n_{\scriptsize\texttt{Max}}\in\{12,16,20,24,28,32\}$ observations are collected.
The usage of group-sequential methods in preclinical experiments has been advocated for in Neumann et al. (2017) and Blotwijk et al. (2022). In a simulation study we compare the approach based on the EUII with three standard group-sequential methods from clinical trials, the Pocock, the O’Brien-Fleming and the Haybittle-Peto stopping rule, as described in the previous section. The first two work with more stringent nominal significance levels at interim to control the Type-I error rate exactly at $\alpha$ in large samples. These results have been extended to also achieve exact T1E control in the $t$-test with small samples (Nikolakopoulos et al., 2016; Rom and McTague, 2020), but for simplicity we use the standard significance levels for large samples available in the R package rpact. The slightly increased T1E rate is automatically taken into account in the computation of EUII, as described in Appendix B, which gives additional details on the design of the simulation study.
Figure 4: The experimental unit information index (first-order) of several group-sequential methods without futility stopping (top) and with futility stopping (bottom).
Figure 5: The experimental unit information index (second-order) of several group-sequential methods without futility stopping (top) and with futility stopping (bottom).
The values of EUII shown in Figure 4 (first-order) and 5 (second-order) are based on simulations for different values of $n_{\scriptsize\texttt{Max}}$, the prior probability $\Pr(H_{1})$ and the true standardized mean difference (SMD). In each Figure, the top panel shows methods which cannot stop early for futility, whereas the bottom plot shows methods which can also stop for futility (including the method from Reinagel). Additional stopping rules for futility has been reported to lead to the saving of resources of up to 30% compared to standard group-sequential methods (Neumann et al., 2017). Here we allow for an additional stop for futility for the group-sequential methods based on the concept of predictive power (PP) (Spiegelhalter et al., 1986; Rufibach et al., 2016). Specifically, we stop the experiment for futility, if the preditive power to achieve significance at the next interim analysis (with four additional observations) is below 30%. The predictive power approach takes into account the uncertainty of the current effect estimate in contrast to standard conditional power calculations (Micheloud and Held, 2022). A threshold on predictive power translates to thresholds on the $p$-values that depend on the interim sample size, as shown in Figure 6. A threshold of 30% leads to $p$-value thresholds between 0.1 and 0.2, whereas smaller thresholds for predictive power (10 or 20%) give larger $p$-values thresholds. We also add futility stopping based on predictive power to $N$-hacking to investigate the effect of a varying $p$-value threshold for futility compared to Reinagel’s approach, where the futility threshold is always 0.1.
Looking at the methods without futility stopping (top panel), we can see that Pocock always has a larger EUII than O’Brien-Fleming, both being better than Haybittle-Peto, similar to Section 4.1. If $\mbox{SMD}=1$ and $n_{\scriptsize\texttt{Max}}$ is relatively large, $N$-hacking (without the additional futility stop of Reinagel’s approach) is even better than Haybittle-Peto and similar to O’Brien-Fleming. The effect of the prior probability $\Pr(H_{1})$ on EUII is relatively small.
The results with the second-order approximation in Figure 5 are qualitatively very similar to the first-order approxmiation shown in Figure 4, with slightly higher values of EUII throughout. The underlying values of $\operatorname{\mathsf{E}}(N_{\bullet})$ and $\mathsf{CV}(N_{\bullet})$ for significant and nonsignificant results are shown in Figure 7 and 8. It can be seen that the expected sample size is smaller for methods that allow for futility stopping, but this is not necessarily the case for the coefficient of variation. Without futility stopping, the coefficient of variation of significant findings can take values up to 0.5, if $n_{\scriptsize\texttt{Max}}$ are large.
Figure 6: Futility bounds on $p$-values based on predictive power.
Comparing the group-sequential methods without futility stopping (top panels) to the approach from Reinagel (dark-orange line in the bottom panels), it is interesting that the approach by Reinagel often performs better in terms of EUII. However, one may argue that this comparison is unfair, because only Reinagel’s “constrained sample augmentation” approach has an additional stop for futility. The bottom panels of Figure 4 and 5 display the resulting EUIIs with futility stopping. The effect of the prior probability on EUII is now somewhat larger, in particular for Haybittle-Peto and O’Brien-Fleming. The EUIIs of the group-sequential methods with additional PP futility stopping increase, with the Pocock or the Haybittle-Peto method now performing best. Pocock has fairly constant values of EUII and is always among the top two methods. Most other methods also have a fairly constant EUII with varying $n_{\scriptsize\texttt{Max}}$, with the exception of Haybittle-Peto, which is increasing from $n_{\scriptsize\texttt{Max}}=16$ to $n_{\scriptsize\texttt{Max}}=32$. As a result, Haybittle-Peto is even better than Pocock if the assumed SMD is small or $n_{\scriptsize\texttt{Max}}$ is large, but can be worse than all other methods if $n_{\scriptsize\texttt{Max}}$ is small. O’Brien-Fleming is better than Reinagel for a small effect size, but for a large effect size it is the other way round. Finally $N$-hacking plus predictive power is very similar to Reinagel’s approach, slightly better for $\mbox{SMD}=1$ and slightly worse otherwise.
In summary, the EUII allowed us to compare designs with different Type-I error rates, power, and expected sample sizes on a scale with direct practical value – the evidentiary value of an additional experimental unit. The second-order approximation gives slightly larger values of EUII, but qualitatively very similar results.
Figure 7: Expected sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
Figure 8: Coefficient of variation of sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
5 Discussion
We have introduced a new measure to quantify the evidentiary value of one experimental unit, the experimental unit information index. Properties in standard statistical test procedures have been described and the corresponding asymptotic values have been derived. Two application have shown the benefits of this measure to compare more complex statistical procedures with efficacy and futility stopping. It would be interesting to investigate more systematically the optimal choice for efficacy and futility stopping in terms of EUII. In the adaptive designs considered in Sections 4.1 and 4.2, we have used a standard overall significance level of 2.5% (one-sided) and 5% (two-sided) to determine the nominal efficacy thresholds of the different adaptive designs considered. Other overall significance values may lead to different (and perhaps larger) values of EUII.
Future work could apply the proposed methodology to dual-criterion designs (Fisch et al., 2015; Roychoudhury et al., 2018; Goodman et al., 2019; Rosenkranz, 2021; Zhao et al., 2023), where success of an experiment requires not only significance, but also relevance of the estimated effect size. Furthermore, Bayesian sequential designs (Gsponer et al., 2014; Gerber and Gsponer, 2016; Pourmohamad and Wang, 2022; Kang et al., 2025) could be considered, where the implied sample size of a prior distribution could be taken into account in the calculation of EUII. Methods to quantify a prior distribution in terms of sample size have been recently proposed (Wiesenfarth and Calderazzo, 2020; Neuenschwander et al., 2020). This aspect is also relevant for Bayesian borrowing methods based on historical data (Gravestock and Held, 2017, 2018). Another possible area of application are more complex experimental designs such as randomized complete block design to incorporate cage effects (Townsend et al., 2025).
There are also interesting possible extensions of the methodology proposed. For example, we may average the power over plausible values of the effect size rather than taking a certain fixed value, as proposed for the pre-experimental rejection ratio (Bayarri et al., 2016). For example, in the absence of precise information on the clinically relevant difference we may use predictive power in the calculation of DOR and EUII. Similarly, we may average the power according to the distribution on the sample sizes of the two groups, because increasing imbalance in sample size leads to increasing loss in power. Finally, in most realistic applications, the EUII must be estimated using simulation. Further research is needed to develop a formula for the Monte Carlo standard error (Koehler et al., 2009) that can quantify the uncertainty associated with the estimate.
Data and Software Availability
Code and data to reproduce our results are openly available at https://doi.org/10.5281/zenodo.17671982. We used the statistical programming language R version 4.4.2 (2024-10-31) for analyses (R Core Team, 2024) along with the xtable (Dahl et al., 2019), biostatUZH (Held et al., 2024), ReplicationSuccess (Held, 2020), ggplot2 (Wickham, 2016), cowplot (Wilke, 2024), ggpubr (Kassambara, 2023), dplyr (Wickham et al., 2022), tidyr (Wickham et al., 2024), rpact (Wassmer and Brannath, 2016), gsdesign (Anderson, 2024), SimDesign (Chalmers and Adkins, 2020), and knitr (Xie, 2024) packages.
Funding
No funding to be declared.
Acknowledgments
We appreciate helpful comments by Tim Friede, Florian Frommlet and Martin Posch.
Conflict of Interest Statement
The authors declare that no conflict of interest exists for all authors.
Author Contributions
Leonhard Held: Conceptualization; Investigation; Writing - original draft; Methodology; Visualization; Writing - review & editing; Formal analysis; Supervision. Fadoua Balabdaoui: Investigation, Methodology, Formal analysis. Samuel Pawel: Writing - review & editing; Writing - original draft; Methodology; Software; Investigation; Visualization; Formal analysis.
Appendix
Appendix A Asymptotic considerations
Here we consider the asymptotic behaviour of the EUII for $n\rightarrow\infty$ for the z-test.
The diagnostic odds ratio can be written as
$\mbox{DOR}=\frac{\mbox{Power Odds}}{\mbox{T1E odds}}=\frac{\mbox{Power}/(1-\mbox{Power})}{\mbox{T1E rate}/(1-\mbox{T1E rate})}\\$
In a standard $z$-test, the term $\mbox{T1E rate}/(1-\mbox{T1E rate})$} in the denominator does not depend on sample size. For example, at the traditional $\alpha=5\%$ significance level, we obtain ${\mbox{T1E rate}/(1-\mbox{T1E rate})}=1/19$.
The term in the numerator, the power odds $\omega=\mbox{Power}/(1-\mbox{Power})$ depend on sample size. In a standard one-sided one-sample $z$-test with sample size $n$, the power is
$\mbox{Power}=\Phi\left(\delta\sqrt{n}-z_{\alpha}\right),$ (23)
here $\Phi(.)$ denotes the standard normal cdf, $\delta>0$ is the assumed standardized effect size of interest and $\alpha$ is the significance level from above. For a two-sample $z$-test with total sample size $n$, the power is
$\mbox{Power}=\Phi\left(\frac{\delta}{2}\sqrt{n}-z_{\alpha}\right).$ (24)
In (23) and (24) we make the common assumption that the probability, that the test rejects the null with a negative effect size in the opposite direction of the true effect $\delta$, can be neglected (Matthews, 2006).
The Power Odds in the one-sample case are hence
$\mbox{Power Odds}=\frac{\Phi(x_{n})}{\Phi(-x_{n})}$
with $x_{n}=a+b\sqrt{n}$ with $a=-z_{\alpha/2}$ and $b=\delta>0$.
We are now interested in the experimental unit information index
$\mbox{EUII}=\mbox{DOR}^{{1}/{n}}$
and want to show that the limit of EUII for $n\rightarrow\infty$ is
$\lim_{n\rightarrow\infty}\mbox{EUII}=\exp\left(\frac{\delta^{2}}{2}\right).$ (25)
It is sufficient to consider the numerator $\left[\frac{\Phi(x_{n})}{\Phi(-x_{n})}\right]^{1/n}$ of EUII because the denominator $\left[\frac{\alpha}{1-\alpha}\right]^{1/n}\rightarrow 1$ for $n\rightarrow\infty$ and any $\alpha<1$.
It remains to be shown that
$\left[\frac{\Phi(x_{n})}{\Phi(-x_{n})}\right]^{1/n}\to\exp(b^{2}/2)$
as $n\to\infty$.
Using $\Phi(-x_{n})=1-\Phi(x_{n})$ we can write
$\left[\frac{\Phi(x_{n})}{\Phi(-x_{n})}\right]^{1/n}=\exp\left(\frac{1}{n}\log\left(\frac{\Phi(x_{n})}{1-\Phi(x_{n})}\right)\right).$
Let $\varphi$ denote the pdf of a standard Gaussian. Using the well known asymptotic behavior of the Mill’s ratio of the Gaussian distribution (Grimmett and Stirzaker, 2001) we have that
$\frac{1-\Phi(t)}{\varphi(t)}\sim\frac{1}{t}$
for $t\to\infty$. This implies that
$\lim_{n\to\infty}\frac{1-\Phi(x_{n})}{\varphi(x_{n})}x_{n}=1.$ (26)
Now,
$\exp\left(\frac{1}{n}\log\left(\frac{\Phi(x_{n})}{1-\Phi(x_{n})}\right)\right)$ $=$ $\exp\left(\frac{1}{n}\log(\Phi(x_{n}))-\frac{1}{n}\log\left(1-\Phi(x_{n})\right)\right)$ (27)
$=$ $\exp\left(\frac{1}{n}\log(\Phi(x_{n}))-\frac{1}{n}\log\left(\frac{1-\Phi(x_{n})}{\varphi(x_{n})}x_{n}\right)-\frac{1}{n}\log(\varphi(x_{n}))+\frac{1}{n}\log(x_{n})\right)$
$\sim$ $\exp\left(-\frac{1}{n}\log(\varphi(x_{n}))\right)$
as a consequence of
$\lim_{n\to\infty}\exp\left(\frac{1}{n}\log(\Phi(x_{n}))\right)=1,$
$\lim_{n\to\infty}\exp\left(\frac{1}{n}\log\left(\frac{1-\Phi(x_{n})}{\varphi(x_{n})}x_{n}\right)\right)=1$
using the limit in (26), and
$\lim_{n\to\infty}\exp\left(\frac{1}{n}\log(x_{n})\right)=1$
since $\log(x_{n})/n=\log(a+b\sqrt{n})/n\sim\log(n)/(2n)$. The last step is to write
$\exp\left(-\frac{1}{n}\log(\varphi(x_{n}))\right)$ $=$ $\exp\left(-\frac{1}{n}\left(-\log(\sqrt{2\pi})-a^{2}/2-ab\sqrt{n}-nb^{2}/2\right)\right)$
$\to$ $\exp(b^{2}/2)$
as $n\to\infty$. Putting this together with (27) completes the proof.
For the two-sample case we replace $b=\delta$ with $b=\delta/2$ and obtain
$\lim_{n\to\infty}\mbox{EUII}_{2}=\exp(b^{2}/8).$
Appendix B Simulation study
Here, we describe additional details on the design of our simulation study, reported in Section 4.2, following the ADEMP reporting structure (Morris et al., 2019; Siepe et al., 2024). This simulation study was not preregistered as it constitutes exploratory research where a new way to quantify the performance of study designs is explored without the intention to give broad recommendations for practitioners (Heinze et al., 2023).
B.1 Aims
The aim of the simulation study is to evaluate the EUII of different sequential design methods, including the sample augmentation approach from Reinagel (2023) and established group-sequential methods with optional stopping for futility based on predictive power.
B.2 Data-generating mechanism
The data-generating mechanism (DGM) closely resembled the DGM from the simulation study by Reinagel (2023). In each simulation repetition, a data set consisting of $n_{\text{Max}}$ observations per group is simulated. The observations in the control group are simulation from a standard normal distribution whereas the observations in the treatment group are simulated from a normal distribution with mean $\delta$ and variance 1. The parameter $\delta$ thus represents the standardized mean difference between treatment and control. Similar to Reinagel, the maximum sample size $n_{\text{Max}}$ is varied from 12 to 32 in steps of 4, representig realistic sample sizes for animal experiments. The standardized mean difference $\delta$ is varied from 0 to 1 to represent no (0), small (0.3), medium (0.5), large (0.8), and very large (1) effect sizes. Both factors are varied in a fully factorial way, leading to $6\times 5=30$ simulation conditions, see Table 4.
Table 4: Simulation factors and their levels.
Factor Levels
Maximum sample size $n_{\text{Max}}\in\{12,16,20,24,28,32\}$
Standardized mean difference $\delta\in\{0,0.3,0.5,0.8,1\}$
B.3 Estimands and other targets
The estimand of interest is the standardized mean difference $\delta$. The target of interest is the null hypothesis $H_{0}\colon\delta=0$.
B.4 Methods
Each simulated data set is analyzed sequentially. Two-sided, two-sample t-tests are conducted using the function stats::t.test(x = ..., y = ..., var.equal = TRUE). The first analysis is performed with 8 observations per group. Subsequent analyses add batches of 4 observations per group until the maximum sample size $n_{\text{Max}}$ is reached.
At each analysis $i=1,\dots,k=(n_{\text{Max}}-8)/4$, the t-test yields a p-value $p_{i}$. Based on $p_{i}$, the null hypothesis $H_{0}$ may or may not be rejected and data collection may or may not stop. If data collection stops at analysis $i$, a total of
$n=8+(i-1)\times 4$
observations per group were used.
For each method, both the rejection decision and the sample size at termination are recorded. The following sequential testing methods are compared:
  • N-hacking: If $p_{1}\leq a_{1}=0.05$ reject $H_{0}$ and stop data collection. Otherwise, continue to the next analysis and repeat this rule with $a_{2}=\ldots,a_{k}=0.05$ until either $H_{0}$ is rejected and data collection stopped, or the final stage is reached without rejection.
  • Reinagel’s sample augmentation design: If $p_{1}\leq 0.05$, reject $H_{0}$ and stop data collection. If $p_{1}>0.1$, stop data collection without rejecting $H_{0}$. If $0.05<p_{1}\leq 0.1$ (the “promising zone”), continue data collection to the next stage. Repeat this rule until data collection stops or the last stage is reached without termination.
  • O’Brien-Fleming: Compute O’Brien-Fleming nominal significance levels $a_{1},\dots,a_{k}$ to maintain overall Type-I error rate of $\alpha=0.05$ using the function rpact::getDesignGroupSequential(...., sided = 2, alpha = 0.05, typeOfDesign = "OF"). At each stage, if $p_{i}\leq a_{i}$, reject $H_{0}$ and stop data collection. Otherwise, continue to the next stage. Repeat this rule until data collection stops or the last stage is reached without termination.
  • Pocock: Compute the Pocock nominal significance level $a_{1}=\ldots=a_{k}=a$ to obtain overall Type-I error rate of $\alpha=0.05$ using the funcion rpact::getDesignGroupSequential(...., sided = 2, alpha = 0.05, typeOfDesign = "P"). At each stage, if $p_{i}\leq a$, reject $H_{0}$ and stop data collection. Otherwise, continue to the next stage. Repeat this rule until data collection stops or the last stage is reached without termination.
  • Haybittle-Peto: Compute Haybittle-Peto nominal significance levels $a_{1}=\ldots=a_{k-1}=0.001,a_{k}=0.05$, which use $a_{i}=0.001$ in the interim analyses and $a_{k}=0.05$ in the final analysis. At each stage, if $p_{i}\leq a_{i}$, reject $H_{0}$ and stop data collection. Otherwise, continue to the next stage. Repeat this rule until data collection stops or the last stage is reached without termination.
Reinagel’s sample augmentation method has the possibility to stop for futility if the $p$-value is greater than 0.1. Because all other methods cannot stop for futility, an additional variant with futility stopping is considered for each of them: In each analysis $i$ where $H_{0}$ is not rejected, the predictive power to reject $H_{0}$ in the next analysis $i+1$ (using the method specific significance level $a_{i+1}$) is calculated with
$\mathrm{PP}_{i+1}=\Phi\left(-t_{1-a_{i+1}/2,2n_{i+1}-2}\sqrt{\frac{1}{1-f}}+z_{i}\sqrt{\frac{f}{1-f}}\right)+\Phi\left(-t_{1-a_{i+1}/2,2n_{i+1}-2}\sqrt{\frac{1}{1-f}}-z_{i}\sqrt{\frac{f}{1-f}}\right)$
where $t_{q,r}$ is the $q\times 100\%$ quantile of the t-distribution with $r$ degrees of freedom, $z_{i}$ is the z-statistic at analysis $i$, and $f=n_{i}/n_{i+1}$ is the ratio of the per-group sample size at analysis $i$ relative to the per-group sample size at analysis $i+1$. Predictive power takes into account the already observed data and their uncertainty (Spiegelhalter et al., 1986; Rufibach et al., 2016; Micheloud and Held, 2022). We specify that if $\mathrm{PP}_{i+1}<30\%$, data collection is stopped without rejection of $H_{0}$, as the probability of still rejecting $H_{0}$ is too low.
B.5 Performance measures
To quantify method performance we first compute the following metrics
  • Rejection rate
    $\widehat{\mathrm{RR}}=\frac{1}{n_{\text{sim}}}\sum_{j=1}^{n_{\text{sim}}}\,\mathbb{1}\left\{\text{Method rejects}\penalty 10000\ H_{0}\colon\delta=0\penalty 10000\ \text{in simulation}\penalty 10000\ j\right\}.$
    which corresponds to the empirical Type-I error rate $\widehat{\mathrm{T1E}}$ in conditions with true $\delta=0$, and the empirical power $\widehat{\mathrm{Pow}}$ otherwise.
  • Average sample size
    $\widehat{\mathrm{E}(N)}=\frac{1}{n_{\scriptsize\text{sim}}}\sum_{j=1}^{n_{\text{sim}}}n_{j}.$
    where $n_{j}$ is the sample size at data collection termination in simulation repetition $j$. Additionally, average sample sizes as in Table 2 were calculated by taking the mean of simulations with only significant or nonsignificant results, respectively, and differentiating between $H_{0}$ conditions (with $\delta=0$) and $H_{1}$ conditions (with $\delta\neq 0$).
We then estimate $\operatorname{\mathsf{E}}(N_{+})$}, $\operatorname{\mathsf{E}}(N_{-})$}, $\operatorname{Var}(N_{+})$} (and likewise $\operatorname{Var}(N_{-})$}) using formulae (20), (21), and (22), respectively, using three different prior probabilities $\Pr(H_{1})\in\{0.01,0.1,0.5\}$. Finally, for each value of $\Pr(H_{1})$, the estimated power, Type-I error rate, and expected sample sizes are combined to compute the first-order EUII with
[…]
and similiarly the second-order $\widetilde{\mbox{EUII}}$ (19), where also $\widehat{\operatorname{Var}}(N_{+})$} and $\widehat{\operatorname{Var}}(N_{-})$} is used.
Each condition was simulated 100’000 times. This ensures a MCSE (Monte Carlo Standard Error) for Type-I error rate and power of at most 0.16%. MCSEs were calculated using the formulae from Siepe et al. (2024), except for the EUII, for which no MCSE formulae is available. However, due to the very large number of repetitions and the minimal variation observed when re-running the simulation with a different seed, Monte Carlo uncertainty reduction was deemed satisfactory.
Convergence was also assessed. The specified DGM always produced valid data sets, all methods always converged without errors, method performance could always be estimated. The simulation study was performed on a computer running Ubuntu 24.04.3 LTS and R version 4.5.2 (2025-10-31). The SimDesign R package was used to organize and run the simulation study (Chalmers and Adkins, 2020). The rpact R package was used to compute group-sequential design bounds (Wassmer and Brannath, 2016).

### v2

@@
-The Experimental Unit Information Index: Balancing Evidentiary Value and Sample Size of Adaptive Designs
+Balancing Evidentiary Value and Sample Size of Adaptive Designs with Application to Animal Experiments
 Abstract:
@@
 Applications to group-sequential designs and a recently proposed adaptive statistical test procedure show the usefulness of the approach when the goal is to maximize the evidentiary value of one experimental unit.
+A reanalysis of 2738 animal experiments with simulated results from (post-hoc) interim analyses illustrates the possible savings in sample size.
 Key Words: Adaptive designs; Diagnostic odds ratio; Likelihood ratio; 3R Principles; Bayesian updating
@@
 Stopping an experiment early ensures that the number of animals included in an experiment does not become too large.
-This reflects the second pillar of the 3R principles “Replace, Reduce, Refine” in animal research (Russell and Burch, 1959), described as
-to use methods for obtaining comparable levels of information from the use of fewer animals in scientific procedures, or for obtaining more information from the same number of animals” (swiss3rcc.org/3rs-for-the-public/)
+These questions are of clear practical relevance in animal research, as exemplified by the recently proposed Stroke Preclinical Assessment Network (SPAN) experimental design (Lyden et al., 2022), calling for more efficient and rigorous designs in preclinical research with stops for both futility and efficacy.
+This also reflects the second pillar of the 3R principles “Replace, Reduce, Refine” in animal research (Russell and Burch, 1959), described as
+to use methods for obtaining comparable levels of information from the use of fewer animals in scientific procedures, or for obtaining more information from the same number of animals” (swiss3rcc.org/3rs-for-the-public/).
 The 3R principles are now widely adopted to improve the design and analysis of animal studies (Reynolds, 2023, 2024a, 2024b; Townsend et al., 2025).
-However, the focus of standard statistical theory on error rates ignores that different adaptive designs may have different (expected) sample sizes.
+However, the focus of standard statistical theory on statistical significance ignores that different adaptive designs may have different (expected) sample sizes.
 Furthermore, a comparison of methods with desired Type-I error rate control with alternative methods that may have more power but also Type-I error rate inflation can be difficult.
-To address the latter issue, optimising the trade-off between Type-I and Type-II error rates has been recently considered Mudge et al.
-(2012); Walley and Grieve (2021); Grieve (2024).
-The goal is here to optimize a weighted sum of error rates based on prior probabilities of the null and the alternative hypothesis and the relative costs of Type-I and Type-II errors.
+To address the latter issue, optimising the trade-off between Type-I and Type-II error rates has been recently proposed (Mudge et al., 2012; Walley and Grieve, 2021; Grieve, 2024).
+The goal is here to optimise a weighted sum of error rates based on prior probabilities of the null and the alternative hypothesis and the relative costs of Type-I and Type-II errors.
 An alternative approach is to adjust power for the corresponding Type-I error rate using the theory of receiver operating characteristic curves (Lloyd, 2005; Cavus et al., 2019).
 However, all these methods do not take sample size into account.
+To put the discussion into context, consider results from 2738 animal experiments in neuroscience and metabolism, collected by Bonapersona et al.
+(2021) through a systematic search of studies that were previously included in meta-analyses published between 2003 and 2019.
+Figure 1 shows the distribution of sample sizes and standardised mean difference effect sizes (Hedges’ $g$).
+Most studies have total sample sizes between 10 and 30 animals but there are some studies with more than 100 animals.
+The majority of standardised mean differences are below 1 in absolute size, but some studies have remarkably large effect sizes.
+This raises the question how many animals could have been saved if the experiments had incorporated a sequential design allowing for interim analyses.
+A related question is which type of group-sequential design (Jennison and Turnbull, 1999) should be used to minimise the expected costs and sample sizes of the experiments while ensuring that the chosen design has sufficiently high evidentiary value as quantified by statistical error rates (Blotwijk et al., 2022).
+Figure 1: Distribution of sample sizes and standardised mean differences in 2738 animal experiments in neuroscience and metabolism (Bonapersona et al., 2021).
+Extreme standardised mean differences with absolute value larger than 5 are not shown ($n=28$).
 In this paper we propose a new measure of statistical information that takes into account Type-I error rate, power, and sample size.
-This measure is related to the diagnostic odds ratio to quantify diagnostic accuracy (Glas et al., 2003), which is widely used in clinical epidemiology and which also has attractive interpretations when applied to statistical hypothesis tests.
+This measure is related to the diagnostic odds ratio to quantify diagnostic accuracy (Glas et al., 2003), which is widely used in clinical epidemiology and has attractive interpretations when applied to statistical hypothesis tests.
 First, it has a frequentist interpretation as the ratio of the rejection odds under the alternative and under the null hypothesis.
 Second, it also has a Bayesian interpretation as the ratio of the posterior odds of the alternative hypothesis given a significant versus nonsignificant results.
-Finally, the diagnostic odds ratio allows for a normalization with respect to the (expected) sample size to obtain a measure of the evidentiary value of one unit/animal/patient, the experimental unit information index (EUII).
+Finally, the diagnostic odds ratio allows for a normalisation with respect to the (expected) sample size to obtain a measure of the evidentiary value of one unit/animal/patient, the experimental unit information index (EUII).
 This article is organized as follows: Section 2 reviews likelihood ratios and diagnostic odds ratios and applies them to statistical hypothesis tests.
-Section 3 then defines the experimental unit information index for fixed sample size and studies its performance in standard statistical test settings.
+Section 3 then defines the experimental unit information index (EUII) for fixed sample size and studies its performance in standard statistical test settings.
 The extension to adaptive tests with different expected sample size of significant and nonsignificant findings is described in Section 4.
 Applications to group-sequential designs (Section 4.1) and “constrained sample augmentation” (Section 4.2), an adaptive test procedure recently proposed by Reinagel (2023), are described.
-We close with some discussion in Section 5.
+In a simulation study we compare the EUII of constrained sample augmentation with group-sequential methods with additional stops for futility based on predictive power.
+Based on our findings, we re-analyse the Bonapersona et al.
+(2021) data set in Section 5 to assess how many animals could have been saved with group-sequential designs, and close with some discussion in Section 6.
 2 Methodology
@@
 In a diagnostic context, $H_{1}$ represents the hypothesis that a patient has the disease whereas $H_{0}$ represents that the patient does not have the disease.
-Likelihood ratios (LRs) are commonly used in diagnostic testing to quantify the (prior to posterior) change in odds $\mathsf{Odds}(H_{1})=\Pr(H_{1})/\Pr(H_{0})$ after having observed a positive or negative test result (Deeks and Altman, 2004):
+Likelihood ratios (LRs) are commonly used to quantify the (prior to posterior) change in odds $\mathsf{Odds}(H_{1})=\Pr(H_{1})/\Pr(H_{0})$ after having observed a positive or negative test result (Deeks and Altman, 2004):
 $\left.\begin{array}[]{c}{\mathsf{Odds}(H_{1}\,|\,\mbox{test positive})}\\ {\mathsf{Odds}(H_{1}\,|\,\mbox{test negative})}\end{array}\right\}$ $=$ $\left\{\begin{array}[]{c}\mbox{LR}^{+}\\ \mbox{LR}^{-}\end{array}\right\}\,\cdot\,{\mathsf{Odds}(H_{1})}.$ (5)
@@
 The diagnostic odds ratio is thus a single indicator of diagnostic test accuracy, the larger, the better.
-We will describe its properties in more detail in the following, when we apply it to statistical hypotheses tests.
+We will describe its properties in more detail in the following, when we apply it to statistical hypothesis tests.
 2.2 The diagnostic odds ratio for hypotheses testing
@@
 These two equations show that likelihood ratios act multiplicatively on the odds of $H_{1}$, transforming prior odds to posterior odds.
-Both likelihood ratios are based on data from $n$ experimental units, so evey unit contributes on average an amount of $({\mbox{LR}^{-}})^{1/n}$ and $({\mbox{LR}^{+}})^{1/n}$, respectively, to the likelihood ratios, because
+Both likelihood ratios are based on data from $n$ experimental units, so every unit contributes on average an amount of $({\mbox{LR}^{-}})^{1/n}$ and $({\mbox{LR}^{+}})^{1/n}$, respectively, to the likelihood ratios, because
 $\mbox{LR}=\underbrace{{\mbox{LR}}^{1/n}\times{\mbox{LR}}^{1/n}\times\ldots\times{\mbox{LR}}^{1/n}}_{\mbox{$n$ times}}.$
@@
 156).
-The increase in DOR is parallelized by an increase in the required sample size $n$ to detect the standardized effect $\delta=0.5$, so the corresponding values of EUII are nearly the same, 1.13 and 1.14, respectively.
+The increase in DOR is parallelized by an increase in the required sample size $n$ to detect the standardised effect $\delta=0.5$, so the corresponding values of EUII are nearly the same, 1.13 and 1.14, respectively.
 The EUII takes its largest value 1.15 if both Type-I and Type-II error have the largest value (5% and 20%, respectively).
@@
 Table 1: Diagnostic odds ratio (DOR) and experimental unit information index (EUII) for different values of power and Type-I error rate.
-Also shown is the required sample size $n$ to detect a standardized effect of $\delta=0.5$ with a one-sample $z$-test and the corresponding value of EUII.
+Also shown is the required sample size $n$ to detect a standardised effect of $\delta=0.5$ with a one-sample $z$-test and the corresponding value of EUII.
 Power (in %) Type-I error rate (in %) DOR $n$ EUII
@@
 90 2.5 351.0 49.6 1.13
-If the design would have been instead based on two equally-sized samples and $\delta$ represents the standardized difference to be detected, then the required total sample size $n=125.6$ at the 5% significance level and 80% power is four times as large (Kirkwood and Sterne, 2003), and we obtain
+If the design would have been instead based on two equally-sized samples and $\delta$ represents the standardised difference to be detected, then the required total sample size $n=125.6$ at the 5% significance level and 80% power is four times as large (Kirkwood and Sterne, 2003), and we obtain
 $\mbox{EUII}_{2}=76^{1/125.6}=1.035.$
 The average evidentiary increase per unit is now 3.5%.
-Note that $\mbox{EUII}_{1}=({\mbox{EUII}_{2}})^{4}$ which means that one unit in a one-sample study contributes as much information as four units in a (perfectly balanced) two-sample trial, if the goal is to detect the same standardized effect.
+Note that $\mbox{EUII}_{1}=({\mbox{EUII}_{2}})^{4}$ which means that one unit in a one-sample study contributes as much information as four units in a (perfectly balanced) two-sample trial, if the goal is to detect the same standardised effect.
 Note that we take $n$ to be the total sample size and not the sample size per group.
@@
 Consider now the standard one-sample one-sided $t$-test to test the null hypothesis $H_{0}\colon\mu=0$ versus the alternative $H_{1}\colon\mu>0$.
-We let the sample size $n$ vary between $2^{3}=8$ and $2^{13}=8192$ and compute the power to detect a standardized effect size $\delta$ at significance level $\alpha=5\%$.
+We let the sample size $n$ vary between $2^{3}=8$ and $2^{13}=8192$ and compute the power to detect a standardised effect size $\delta$ at significance level $\alpha=5\%$.
 Figure 3 shows how the resulting EUII varies with $\delta$ and $n$.
@@
 Figure 3: The experimental unit information index for the standard one-sample one-sided $t$-and $z$-test as a function of sample size and effect size.
-The asymptotic value $\exp(\delta^{2}/2)$ for the different values of the standardized effect size $\delta$ is indicated by the dotted lines.
+The asymptotic value $\exp(\delta^{2}/2)$ for the different values of the standardised effect size $\delta$ is indicated by the dotted lines.
 Note that the power also increases with increasing $n$, but the EUII index adjusts for the sample size $n$ and is able to compensate for this increase.
@@
 In the following, we denote a random sample size by $N$ to distinguish it from a fixed sample size $n$.
-An adaptive trial can then be characterized by the pair of the sample size $N$ at which the trial stops and the indicator $S$ of whether the results were (non)significant.
+An adaptive trial can then be characterised by the pair of the sample size $N$ at which the trial stops and the indicator $S$ of whether the results were (non)significant.
 The expected sample sizes for significant (++) and nonsignificant (-) are then then defined by the conditional expectations $\operatorname{\mathsf{E}}(N_{+})=\operatorname{\mathsf{E}}(N\,|\,S=\text{significant})$ and $\operatorname{\mathsf{E}}(N_{-})=\operatorname{\mathsf{E}}(N\,|\,S=\text{nonsignificant})$, respectively.
@@
 The rationale for the generalization (18) is that the computation of posterior predictive odds with likelihood ratios is either based on significant (equation (8)) or nonsignificant findings (equation (9)).
-It is therefore natural to normalize the numerator $\mbox{LR}^{+}$ in (18) with the expected sample size of significant findings $\operatorname{\mathsf{E}}(N_{+})$ and the denominator with the expected sample size of nonsignificant findings $\operatorname{\mathsf{E}}(N_{-})$.
+It is therefore natural to normalise the numerator $\mbox{LR}^{+}$ in (18) with the expected sample size of significant findings $\operatorname{\mathsf{E}}(N_{+})$ and the denominator with the expected sample size of nonsignificant findings $\operatorname{\mathsf{E}}(N_{-})$.
 For designs with fixed sample size $n=\operatorname{\mathsf{E}}(N_{+})=\operatorname{\mathsf{E}}(N_{-})$, this generalization reduces correctly to (14).
@@
 Note that Pocock and O’Brien-Fleming exactly control the Type-I error rate at $\alpha=2.5\%$ whereas the one-sided Haybittle-Peto method (Blenkinsop et al., 2019) (with nominal thresholds $a_{1}=a_{2}=a_{3}=0.0005$ and $a_{4}=0.025$) has a slightly increased overall T1E rate of 2.54%.
-The maximum sample size $n_{\scriptsize\texttt{Max}}$ (last column of Table 3) from Pocock and O’Brien-Fleming is larger than the corresponding sample size 50 under a fixed design with the same Type-I error rate and power of 90% (corresponding to a standardized effect size of $\delta=0.46$).
+The maximum sample size $n_{\scriptsize\texttt{Max}}$ (last column of Table 3) from Pocock and O’Brien-Fleming is larger than the corresponding sample size 50 under a fixed design with the same Type-I error rate and power of 90% (corresponding to a standardised effect size of $\delta=0.46$).
 Without futility stopping, this is also the sample size for any nonsignificant result, both under the null and the alternative: $\operatorname{\mathsf{E}}(N_{0-})=\operatorname{\mathsf{E}}(N_{1-})=\operatorname{\mathsf{E}}(N_{-})=n_{\scriptsize\texttt{Max}}$.
@@
 Although the power and Type-I error rate of Haybittle-Peto does not correspond in this case to the other methods, the EUII enables a meaningful comparison.
-Figure 3 displays the EUII for the different methods (top), relative to the fixed design (middle), and the maximum sample size and expected sample size of significant findings (bottom) for values of the standardized effect size $\delta$ between 0.1 and 0.8.
+Figure 4 displays the EUII for the different methods (top), relative to the fixed design (middle), and the maximum sample size and expected sample size of significant findings (bottom) for values of the standardised effect size $\delta$ between 0.1 and 0.8.
 As expected, the EUII increases with increasing $\delta$ but also depends on the assumed prior probability of $H_{1}$, chosen to be either 0.01, 0.1 or 0.5.
@@
 These results indicate that Pocock should be the preferred group-sequential method if the goal is to maximize the evidentiary value of one experimental unit.
-Figure 3 also shows the expected sample size of significant findings under the two hypotheses, for the different values of the standardized effect size.
+Figure 4 also shows the expected sample size of significant findings under the two hypotheses, for the different values of the standardised effect size.
 With the exception of Pocock, the expected sample size of significant findings is always larger under $H_{0}$ than under $H_{1}$.
-Only for Pocock it can be the other way round if the standardized effect size is less than $\delta=0.46$, which corresponds to the power being below 90%.
+Only for Pocock it can be the other way round if the standardised effect size is less than $\delta=0.46$, which corresponds to the power being below 90%.
 This is a consequence of the fact that, under $H_{0}$, a significant finding with Pocock’s method is more likely to be obtained at the first interim analysis and less at the second and so forth.
@@
 This peculiar feature is caused by the constant and relatively large nominal significance levels (see Table 3) and has been one of the criticisms of the Pocock method in clinical trials, where – under the $H_{0}$ – it is less difficult that a trial is stopped early than for competing methods (Simon, 1994; Matthews, 2006).
-However, this feature has been described as an adventage in animal trials where the goal is to minimize animal numbers (Blotwijk et al., 2022) and our analysis confirms that smaller expected sample sizes under $H_{0}$ lead to a larger evidentiary value per animal of the Pocock method compared to the other ones.
+However, this feature has been described as an advantage in animal trials where the goal is to minimize animal numbers (Blotwijk et al., 2022) and our analysis confirms that smaller expected sample sizes under $H_{0}$ lead to a larger evidentiary value per animal of the Pocock method compared to the other ones.
 For example, the EUII is smaller for Haybittle-Peto, where the nominal significance levels (with the exception of the last one) are also constant, but smaller.
@@
 Figure 6: The experimental unit information index (second-order) of several group-sequential methods without futility stopping (top) and with futility stopping (bottom).
-The values of EUII shown in Figure 4 (first-order) and 5 (second-order) are based on simulations for different values of $n_{\scriptsize\texttt{Max}}$, the prior probability $\Pr(H_{1})$ and the true standardized mean difference (SMD).
+The values of EUII shown in Figure 5 (first-order) and 6 (second-order) are based on simulations for different values of $n_{\scriptsize\texttt{Max}}$, the prior probability $\Pr(H_{1})$ and the true standardised mean difference (SMD).
 In each Figure, the top panel shows methods which cannot stop early for futility, whereas the bottom plot shows methods which can also stop for futility (including the method from Reinagel).
@@
 Here we allow for an additional stop for futility for the group-sequential methods based on the concept of predictive power (PP) (Spiegelhalter et al., 1986; Rufibach et al., 2016).
-Specifically, we stop the experiment for futility, if the preditive power to achieve significance at the next interim analysis (with four additional observations) is below 30%.
+Specifically, we stop the experiment for futility, if the predictive power to achieve significance at the next interim analysis (with four additional observations) is below 30%.
 The predictive power approach takes into account the uncertainty of the current effect estimate in contrast to standard conditional power calculations (Micheloud and Held, 2022).
@@
 The effect of the prior probability $\Pr(H_{1})$ on EUII is relatively small.
-The results with the second-order approximation in Figure 5 are qualitatively very similar to the first-order approxmiation shown in Figure 4, with slightly higher values of EUII throughout.
-The underlying values of $\operatorname{\mathsf{E}}(N_{\bullet})$ and $\mathsf{CV}(N_{\bullet})$ for significant and nonsignificant results are shown in Figure 7 and 8.
+The results with the second-order approximation in Figure 6 are qualitatively very similar to the first-order approximation shown in Figure 5, with slightly higher values of EUII throughout.
+The underlying values of $\operatorname{\mathsf{E}}(N_{\bullet})$ and $\mathsf{CV}(N_{\bullet})$ for significant and nonsignificant results are shown in Figures B1 and B2 in Appendix B.
 It can be seen that the expected sample size is smaller for methods that allow for futility stopping, but this is not necessarily the case for the coefficient of variation.
@@
 However, one may argue that this comparison is unfair, because only Reinagel’s “constrained sample augmentation” approach has an additional stop for futility.
-The bottom panels of Figure 4 and 5 display the resulting EUIIs with futility stopping.
+The bottom panels of Figure 5 and 6 display the resulting EUIIs with futility stopping.
 The effect of the prior probability on EUII is now somewhat larger, in particular for Haybittle-Peto and O’Brien-Fleming.
@@
 The second-order approximation gives slightly larger values of EUII, but qualitatively very similar results.
-Figure 7: Expected sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
-Figure 8: Coefficient of variation of sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
+5 Application: How many animals could be saved with group-sequential designs?
+We will now revisit the 2738 animal experiments in neuroscience and metabolism from Bonapersona et al.
+(2021).
+We will assess how many animals could have been saved had each experiment also featured an interim analysis.
+We will also investigate the rejection rates of different group-sequential designs with and without additional stops for futility.
+Assume that the interim analysis takes places after the data from $n_{1}$ animals have been collected, whereas the final analysis takes place after $n_{2}>n_{1}$ animals.
+The corresponding z-values have the marginal distributions $Z_{i}\sim\mathrm{N}(\delta\sqrt{n_{i}},1)$ for $i\in\{1,2\}$, while their correlation is $\mathrm{Cor}(Z_{1},Z_{2})=\sqrt{n_{1}/n_{2}}$ (Jennison and Turnbull, 1999).
+For one-group comparisons, $n_{1}$ and $n_{2}$ are the actual sample sizes at each stage, but for two-group comparisons, as in the Bonapersona et al.
+(2021) data, they are the effective sample sizes given by half the harmonic mean of the sample sizes in the control and treatment groups.
+In the data from Bonapersona et al.
+(2021), we only know the observed value of $z_{2}$ at the final analysis, but can derive the conditional distribution of $Z_{1}$ given the observed $z_{2}$, which is
+$Z_{1}\mid Z_{2}=z_{2}\sim\mathrm{N}\left(z_{2}\sqrt{n_{1}/n_{2}},1-n_{1}/n_{2}\right).$ (23)
+Importantly, the standardised effect $\delta$ cancels out, so the conditional distribution of $Z_{1}$ given $Z_{2}=z_{2}$ is the same regardless of whether data are generated under the null $H_{0}\colon\delta=0$ or the alternative $H_{1}\colon\delta\neq 0$.
+We will now use this result and simulate interim analyses to assess the effect of early stopping.
+Table 4: Mean sample size and rejection rate when introducing an additional interim analyses in 2’738 animal experiments in neuroscience and metabolism (Bonapersona et al., 2021).
+Values represent medians with 2.5% and 97.5% quantiles from 10’000 simulated interim analyses.
+Methods with futility stopping halt the study if the predictive power for rejecting $H_{0}$ in the final analysis is less than 10%.
+Method Mean $n$ $H_{0}$ rejections (%) Interim efficacy stops (%) Interim futility stops (%) Animals saved
+No interim analysis 22.5 27.8 0 0 0
+Haybittle-Peto 21.2 (21.1, 21.3) 28.3 (28, 28.5) 11.8 (11, 12.5) 0 (0, 0) 3’804 (3’515, 4’099)
+O’Brien-Fleming 21.4 (21.3, 21.5) 27.6 (27.4, 27.8) 9.8 (9.1, 10.4) 0 (0, 0) 3’152 (2’891, 3’426)
+Pocock 20.6 (20.5, 20.7) 25.9 (25.3, 26.4) 16.7 (15.9, 17.6) 0 (0, 0) 5’384 (5’043, 5’738)
+Haybittle-Peto + futility 18.2 (18, 18.4) 27.6 (27.2, 28) 11.8 (11, 12.5) 28.7 (27.2, 30.2) 11’858 (11’263, 12’470)
+O’Brien-Fleming + futility 18.4 (18.2, 18.6) 26.9 (26.6, 27.2) 9.8 (9.1, 10.4) 29.4 (27.9, 31) 11’423 (10’832, 12’024)
+Pocock + futility 16.7 (16.5, 17) 25.3 (24.7, 25.9) 16.7 (15.9, 17.6) 37.1 (35.6, 38.7) 15’905 (15’286, 16’541)
+Table 4 shows results from simulating 10’000 interim $z$-statistics $Z_{1}$ from (23) for each observed final $z$-statistic $z_{2}$ in the Bonapersona et al.
+(2021) data set.
+The interim analyses were conducted halfway through the experiment, where the the number of animals in control and treatment groups has been divided by two and rounded to the next integer, if non-integer values were obtained.
+We considered Haybittle-Peto, O’Brien-Fleming, and Pocock stopping rules, both with and without possibility for futility stopping.
+Following the recommendation by Matthews (2006), we use a nominal interim significance threshold of 0.01 (rather than 0.001) for the Haybittle-Peto method with just one interim analysis.
+This results in a further inflated Type-I error rate, whereas the other two group-sequential methods control the Type-I error rate at the chosen two-sided $\alpha=5\%$ level.
+We have used a futility stopping threshold of 10% for the predictive power at the end of the experiment.
+Without futility stopping, the mean sample size is reduced by around 1-2 animals per experiment compared to no interim analysis.
+However, further reductions can be achieved with additional futility stopping.
+The Pocock design with futility stopping saves nearly 6 animals per experiment, totalling to around 16’000 animals.
+This is accompanied by a drop in $H_{0}$ rejection rates from 27.8% to 25.3%.
+It should be noted that reduced rejection rates of adaptive designs are not necessarily a disadvantage.
+In the Bonapersona et al.
+(2021) data set, the ground truth is unknown, and as such we cannot know whether a $H_{0}$ rejection is a true or false positive.
+Note that the Haybittle-Peto design has even higher rejection rates than without an interim analysis because of its inflated Type-I error rate.
+Interestingly, 37.1% of experiments stop for futility with the Pocock design, whereas this number is below 30% for the other two group-sequential designs due to their larger nominal significance thresholds at the final analysis.
+Stops for efficacy are also more common with the Pocock design, where 1 in 6 experiments could have been stopped early for efficacy, compared to 1 in 8 with the Haybittle-Peto design and 1 in 10 with O’Brien-Fleming.
+In summary, using group-sequential designs that allow for futility stops could save a large number of resources and animals.
+In particular, the Pocock method combined with futility stopping results in the largest savings in sample size, with only modest reductions in $H_{0}$ rejections.
+This is consistent with the findings from our simulation study described in Section 4.2, where Pocock emerged as the best method in terms of EUII, admittedly in a setting with several (rather than one) interim analyses.
 6 Discussion
@@
 Properties in standard statistical test procedures have been described and the corresponding asymptotic values have been derived.
-Two application have shown the benefits of this measure to compare more complex statistical procedures with efficacy and futility stopping.
+Two applications have shown the benefits of this measure to compare more complex statistical procedures with efficacy and futility stopping.
 It would be interesting to investigate more systematically the optimal choice for efficacy and futility stopping in terms of EUII.
@@
 Data and Software Availability
+The CC-By 4.0 licensed data set from Bonapersona et al.
+(2021) was downloaded from https://doi.org/10.17605/OSF.IO/WVS7M.
 Code and data to reproduce our results are openly available at https://doi.org/10.5281/zenodo.17671982.
@@
 Acknowledgments
-We appreciate helpful comments by Tim Friede, Florian Frommlet and Martin Posch.
+We appreciate helpful comments by Tim Friede, Florian Frommlet and Martin Posch. We thank Bonapersona et al. (2021) for openly sharing their data.
 Conflict of Interest Statement
@@
 $\mbox{Power}=\Phi\left(\delta\sqrt{n}-z_{\alpha}\right),$ (24)
-here $\Phi(.)$ denotes the standard normal cdf, $\delta>0$ is the assumed standardized effect size of interest and $\alpha$ is the significance level from above.
+here $\Phi(.)$ denotes the standard normal cdf, $\delta>0$ is the assumed standardised effect size of interest and $\alpha$ is the significance level from above.
 For a two-sample $z$-test with total sample size $n$, the power is
@@
 The observations in the control group are simulation from a standard normal distribution whereas the observations in the treatment group are simulated from a normal distribution with mean $\delta$ and variance 1.
-The parameter $\delta$ thus represents the standardized mean difference between treatment and control.
+The parameter $\delta$ thus represents the standardised mean difference between treatment and control.
 Similar to Reinagel, the maximum sample size $n_{\text{Max}}$ is varied from 12 to 32 in steps of 4, representig realistic sample sizes for animal experiments.
-The standardized mean difference $\delta$ is varied from 0 to 1 to represent no (0), small (0.3), medium (0.5), large (0.8), and very large (1) effect sizes.
+The standardised mean difference $\delta$ is varied from 0 to 1 to represent no (0), small (0.3), medium (0.5), large (0.8), and very large (1) effect sizes.
 Both factors are varied in a fully factorial way, leading to $6\times 5=30$ simulation conditions, see Table 5.
@@
 Maximum sample size $n_{\text{Max}}\in\{12,16,20,24,28,32\}$
-Standardized mean difference $\delta\in\{0,0.3,0.5,0.8,1\}$
+Standardised mean difference $\delta\in\{0,0.3,0.5,0.8,1\}$
 B.3 Estimands and other targets
-The estimand of interest is the standardized mean difference $\delta$. The target of interest is the null hypothesis $H_{0}\colon\delta=0$.
+The estimand of interest is the standardised mean difference $\delta$. The target of interest is the null hypothesis $H_{0}\colon\delta=0$.
 B.4 Methods
@@
 Because all other methods cannot stop for futility, an additional variant with futility stopping is considered for each of them: In each analysis $i$ where $H_{0}$ is not rejected, the predictive power to reject $H_{0}$ in the next analysis $i+1$ (using the method specific significance level $a_{i+1}$) is calculated with
-$\mathrm{PP}_{i+1}=\Phi\left(-t_{1-a_{i+1}/2,2n_{i+1}-2}\sqrt{\frac{1}{1-f}}+z_{i}\sqrt{\frac{f}{1-f}}\right)+\Phi\left(-t_{1-a_{i+1}/2,2n_{i+1}-2}\sqrt{\frac{1}{1-f}}-z_{i}\sqrt{\frac{f}{1-f}}\right)$
+$\mathrm{PP}_{i+1}=\Phi\left(-\frac{t_{1-a_{i+1}/2,2n_{i+1}-2}}{\sqrt{1/f-1}}+\frac{z_{i}}{\sqrt{1-f}}\right)+\Phi\left(-\frac{t_{1-a_{i+1}/2,2n_{i+1}-2}}{\sqrt{1/f-1}}-\frac{z_{i}}{\sqrt{1-f}}\right)$
 where $t_{q,r}$ is the $q\times 100\%$ quantile of the t-distribution with $r$ degrees of freedom, $z_{i}$ is the z-statistic at analysis $i$, and $f=n_{i}/n_{i+1}$ is the ratio of the per-group sample size at analysis $i$ relative to the per-group sample size at analysis $i+1$.
@@
 The rpact R package was used to compute group-sequential design bounds (Wassmer and Brannath, 2016).
+B.6 Additional results
+Figures B1 and B2 shows the means and coefficient of variations of sample sizes with significant and nonsignificant results.
+Figure B1: Expected sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
+Figure B2: Coefficient of variation of sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.